Wesley Company manufactures and sells a single product. The company’s sales and expenses for last quarter follow:


Total Per Unit
  Sales $ 750,000 $ 30
  Less: Variable expenses 525,000 21
  Contribution margin 225,000 $ 9
  Less: Fixed expenses 180,000
  Net operating income $ 45,000


Required:
1. What is the quarterly break-even point in units sold and in sales dollars?

      

2. Without resorting to computations, calculate the total contribution margin at the break-even point.

      

3. How many units would have to be sold each quarter to earn a target profit of $36,900? Use the formula method.

      

4.

Refer to the original data. Compute the company’s margin of safety for the quarter in both dollar and percentage terms. (Round "Percentage" answer to 1 decimal place, (i.e., 0.123 should be considered as 12.3%).)

      

5.

What is the company’s CM ratio? If quarterly sales increase by $22,500 and there is no change in fixed expenses, by how much would you expect quarterly net operating income to increase? (Do not prepare an income statement; use the CM ratio to compute your answer.)
Answer #1

Answers

  • [1]

A

Fixed expenses

$180,000

B

Contribution margin per unit

$9

C = A/B

Break even point in units

20000

D = C x $ 30

Break even point in sales $

$600,000

  • [2]

Total contribution margin at break even point = Fixed expenses = $ 180,000

  • [3]

Units required to be sold = (Target profit + Fixed expenses) / Contribution margin per unit
= ($36900 + 180000) / $ 9
= 216900 / 9
= 24,100 units

  • [4]

A

Total Sales

$750,000

B

Break even point in sales $

$600,000

C = A - B

Margin of Safety $

$150,000

D = C/A) x 100

Margin of Safety %

20%

  • [5]

A

Contribution margin per unit

$9

B

Sale price per unit

$30

C = (A/B) x 100

CM Ratio

30%

D

Increase in sales

$22,500

E = C x D

Increase in Net Operating Income

$6,750